1) Given the reaction: 2CO + O​2 ​→ 2CO​2

What is the minimum number of moles of O​2 ​required to produce one mole of CO​2​?


2) In the reaction Zn + 2HCl ​ ​→ ZnCl​2 ​+ H​2​,

How many moles of hydrogen will be formed when 4 moles of HCl are consumed?


3) Given the reaction: 4Al + 3O​2 ​→ 2Al​2​O​3

How many moles of Al​2​O​3 ​will be formed when 27 grams of Al reacts completely with O​2​?


4) In the reaction N​2 ​+ 3H​2 ​→ 2NH​3​,

How many grams of hydrogen are needed to produce exactly 1 mole of ammonia?


5) Given the balanced equation: NaOH + HCl ​ ​→ NaCl + H​2​O 

 What is the total number of grams of H​2​O produced when 116 grams of the product, NaCl, is formed


6) Given the reaction: 4NH​3(g) ​+ 5O​2(g) ​→ 4NO​(g) ​+ 6H​2​O​(g)

At constant pressure, how many liters of O​2(g) ​would be required to produce 40. liters of NO​(g)​?


1
Expert's answer
2021-01-15T06:35:16-0500

1) 1/2 moles.

2) 4/2 = 2 moles.

3) 1 x 2/4= 1/2 moles.

4) 6/2 = 3 g.

5) 116x18/58.5= 36 g.

6) 50 L.